A Selecting output in perfect competition. Expected length of no more than 1 page

IVA. For managers operating in perfect competition, explain and justify the rule MR = MC for selecting output to maximize economic profit. Explain the relationship between the P = MC and the rule MR = MC to select the output level. That is, compare and contrast the two rules in a perfectly competitive market.

For managers to maximize economic profit: select an output level where MR = MC

For managers to maximize total surplus:         select an output level where P     = MC

Explanations should be written in an essay format. If you use to use graphs to justify or demonstrate your point, just copy an example.
Answer #1

In perfect competition, each firm faces a horizontal demand curve which means that marginal revenue curve(MR) is also horizontal and parallel to x-axis. Now suppose total revenue of the firm is TR and total cost is TC. We know that dTR/dQ and dTC/dQ is equal to MR and MC. Now, we have profit for tge firm equal to,

Profit = TR - TC

To maximise this profit we have to differentiate the the profit function with respect to quantity and equate it to zero, this will give us the profit maximising quantity,

dProfit/dQ = dTR/dQ - dTC/dQ = MR - MC = 0

=> MR = MC,

This gives us the profit maximizing quantity Q.

Now, we know that the firm faces a horizontal MR curve, so, corresponding to the profit maximizing quantity, the price will be equal to MR and thus, MC. At this price the total surplus would be maximized because at P = MC, each consumer is able to buy the product who is willing to pay atleast the cost of producing an extra unit. Thus, total surplus is maximized.

*Please don’t forget to hit the thumbs up button, if you find the answer helpful.

Thank You